# Principal Components for Neural Network Initialization

### v1

@@ III-C Theoretical analysis
PCA transforms the data into uncorrelated features. Therefore, the Hessian matrix (second derivative of the loss) becomes more well-conditioned after PCA, making optimization more stable.
Theorem 1
Consider training a single-layer neural network for a linear regression problem, where we have the input matrix $X$ in $\mathbb{R}^{n\times d}$, the label $\hat{Y}$ in $\mathbb{R}^{n\times 1}$, and we are using the weight matrix $W$ in $\mathbb{R}^{d\times 1}$. Then, PCA removes small eigenvalues of the Hessian matrix of the MSE loss function, leading to faster convergence.
Proof 1
It is known that the steepest descent method for nonlinear optimization is highly sensitive to scaling and the problem’s condition number and that convergence can be significantly slow when the condition number of the Hessian matrix is large [11].
For a linear regression problem with single-layer neural network, input matrix $X$, the label $\hat{Y}$, and the weight matrix $W$, the Mean Squared Error (MSE) loss function is
$L(W)=\frac{1}{2}\|XW-Y\|^{2}.$
Therefore, the gradient of the loss function with respect to $W$ is:
$\nabla L(W)=X^{T}(XW-Y),$
and the second derivative (Hessian) is
$H=\nabla^{2}L(W)=X^{T}X$
Thus, the Hessian depends on the structure of $X^{T}X$. Its eigenvalues determine how efficiently gradient descent converges. The condition number of the Hessian matrix is:
$\kappa(H)=\frac{\lambda_{\max}}{\lambda_{\min}}$
where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalues of $X^{T}X$.
When input features are highly correlated, some eigenvalues of $X^{T}X$ become very small, making $\kappa(H)$ large and slowing optimization.
PCA transforms the input $X$ into uncorrelated principal components $Z$ using an orthonormal transformation matrix $V$:
$Z=XV$
Since $V$ is an orthonormal matrix (i.e., $V^{T}V=I$), the covariance of $Z$ is:
$\text{Cov}(Z)=V^{T}X^{T}XV$
By construction, PCA diagonalizes $X^{T}X$, meaning the Hessian in the transformed space becomes:
$H^{\prime}=Z^{T}Z=\Lambda$
where $\Lambda$ is a diagonal matrix containing the eigenvalues of $X^{T}X$.
Since PCA removes the smallest eigenvalues (low-variance directions), the condition number of $H^{\prime}$ is lower, leading to faster convergence.
IV Experiments
In this section, we will illustrate the performance of PCsInit and its variations compared to training neural network based on principal components (PCA-NN), and the performance of neural network without PCA (NN).
IV-A Experiment Settings
To illustrate the effectiveness of the proposed initialization strategies, we conduct various experiments on various datasets. The descriptions of the datasets used are as in Table I.
For each experiment, we choose the number of principal components so that a minimum of 95% of variance is retained, as commonly used in various research articles [12, 13, 14, 15]. The width of the network used for each dataset has a width equal to the number of principal components retained. All models are trained with a total of 200 epochs using Cross Entropy Loss and Adam optimizer with a learning rate of $0.01$. The ReLU activation function is applied to all layers except the last layer. The number of layers used for all datasets is 5.
In PCsInit and its variations, the first layer is initialized with the principal components, and then it is frozen for the first 30 epochs. After that, we unfreeze this first layer and train the whole model for 170 more epochs. To facilitate a fair comparison, for all strategies, all layers are initialized with the same weights, except for the first layer, since PCsInit and its variants are initialized based on principal components, while PC-NN and NN are not. We compared each dataset with three different initialized techniques: He initialization, Xavier initialization, and Orthogonal initialization, and report the performance.
The experiments run on a CPU of AMD Ryzen 3 3100 4-Core Processor, installed RAM 16.0 GB. The codes for the experiments is available at github.com/pthnhan/pcsinit.
TABLE I: Descriptions of datasets used in the experiments
@@ IV-A Experiment Settings
Heart $2$ $44$ $267$
Ionosphere $2$ $32$ $351$
Micromass $2$ $1087$ $360$
Parkinson $2$ $754$ $756$
The five datasets we evaluated are all of which involve binary classification and are categorized into two primary groups based on the number of features. The first group includes datasets with a smaller number of features such as Heart and Ionosphere. The second group consists of datasets with a larger number of features: MicroMass and Parkinson.
IV-B Results and Analysis
Our experiments cover a diverse set of datasets, ranging from datasets with few features to those where the number of features exceeds the number of samples (Micromass dataset). The experiments on these datasets illustrate the outperformance results of PCsInit, PCsInit-Act, PCsInit-sub, and PCA-NN compared to NN. Moreover, PCsInit and its variations also show excellent results in various data sets under three initialization techniques with better results compared to PCA-NN in both testing accuracy and convergence speed.
The experimental results demonstrate that PCsInit variations and PCA-NN outperform the standard NN initialization across most datasets and initialization techniques. In many cases, such as the Heart dataset, PCsInit consistently outperforms PCA-NN across all initialization techniques (figures 3, 3, 3), highlighting its adaptability and efficiency in optimizing the network’s performance.
In some cases, PCsInit shows performance comparable to PCA-NN, particularly evident in the ionosphere dataset, where both methods achieve similar levels of accuracy and training loss stability. On the other hand, PCA-NN shows its strength on the micromass dataset, achieving higher accuracy compared to PCsInit. Notably, PCsInit-Sub on this dataset exhibits competitive performance with PCA-NN, demonstrating that it can achieve similar results while reducing computational costs. These findings underscore that while both PCsInit and PCA-NN provide substantial improvements over NN, the variations of PCsInit, such as PCsInit-Sub and PCsInit-Act, offer additional flexibility and adaptability, enabling robust performance across a broader range of scenarios.
The results demonstrate that PCsInit-Sub achieves competitive performance compared to the full PCsInit method, while significantly reducing computational overhead. Across all datasets, PCsInit-Sub consistently delivers similar levels of accuracy and training loss stability, closely mirroring the trends observed in PCsInit. For example, on the Heart dataset, PCsInit-Sub performs better under He initialization (figure 3), slightly worse under Orthogonal initialization (figure 3), and achieves the same performance under Xavier initialization (3). On the Ionosphere dataset, PCsInit-Sub demonstrates mostly the same performance across all initialization techniques (figures 6, 6, and 6), maintaining its robustness. Additionally, on the Micromass dataset, PCsInit-Sub exhibits more stable and better performance under He initialization (figure 9), while on the Parkinson dataset, it outperforms PCsInit under Xavier initialization (figure 12). This ability to maintain high performance, stability, and reduced computational speed as indicated in figures 14 and 14 highlights the efficiency of PCsInit-Sub, making it an appealing alternative for scenarios with large datasets or limited resources.
Next, the results also show that PCsInit-Act many times can be a more stable and competitive variation compared to the standard PCsInit, with slight improvements observed on the Ionosphere and Micromass datasets across all initialization techniques (figures 6, 6, 6, 9, 9, and 9). This highlights its ability to adapt effectively to varying data structures while maintaining consistent performance. Notably, PCsInit-Act demonstrates exceptional results under the Xavier initialization technique for other layers, emerging as the best-performing approach on three out of the four datasets: Heart, Ionosphere, and Micromass (figures 3, 6, and 9). This consistent superiority under Xavier initialization underscores the robustness of PCsInit-Act in leveraging the initialization method to enhance training stability and accuracy. These findings emphasize the effectiveness of PCsInit-Act, particularly in scenarios requiring both competitive performance and stability.
For datasets with a larger number of features, such as the Micromass dataset where the number of features exceeds the number of samples, the challenges of training are more pronounced. While PCsInit does not achieve exceptional results across all initialization techniques, it still outperforms the standard NN, demonstrating its robustness in handling high-dimensional data. Furthermore, its variations, PCsInit-Act and PCsInit-Sub, exhibit superior performance, offering attractive alternative approaches with enhanced stability and competitive accuracy. These results highlight the adaptability and effectiveness of the proposed methodology, particularly in addressing the complexities of high-dimensional datasets.
Figure 1: Heart dataset with 5 layers, He initialization
@@ IV-B Results and Analysis
Figure 2: Heart dataset with 5 layers, Xavier initialization
Figure 3: Heart dataset with 5 layers, Orthogonal initialization
Figure 4: Ionosphere dataset with 5 layers, He initialization
@@
Figure 5: Ionosphere dataset with 5 layers, Xavier initialization
Figure 6: Ionosphere dataset with 5 layers, Orthogonal initialization
Figure 7: Micromass dataset with 5 layers, He initialization
@@
Figure 8: Micromass dataset with 5 layers, Xavier initialization
Figure 9: Micromass dataset with 5 layers, Orthogonal initialization
Figure 10: Parkinson dataset with 5 layers, He initialization
@@
Figure 11: Parkinson dataset with 5 layers, Xavier initialization
Figure 12: Parkinson dataset with 5 layers, Orthogonal initialization
Figure 13: Micromass dataset with 5 layers, He initialization
Figure 14: Parkinson dataset with 5 layers, He initialization
V Conclusion
In conclusion, in this work, we propose PCsInit, which initializes the first layer of a neural network with the principal components obtained from the input data and illustrates its advantages compared to training a neural network on principal components. In general, PCsInit and its variations (PCsInit-Sub and PCsInit-Act) not only provide a more direct and easy way to explain the decision of a model but also outperform PCA-NN and NN across various datasets and initialization techniques. We also show theoretically and experimentally that PCsInit and its variants converge faster than NN.
@@ V Conclusion
In the future, we will examine the performance of the proposed approaches under various neural network architectures and various types of data (noisy, imbalanced). In addition, we will explore the potential usage of Kernel PCA instead of PCA for initializing neural networks.

### v2

@@ 2.2.1 Convergence analysis
PCA transforms the data into uncorrelated features. Therefore, the Hessian matrix (second derivative of the loss) becomes better-conditioned after PCA, making optimization more stable. While rigorous analysis for the whole neural network is challenging, we offer a simplifed case scenario to illustrate the performance of PCsInit in a simplified settings:
Theorem 2.1
Consider a single-layer linear neural network regression model with input data $X\in\mathbb{R}^{d\times n}$ (where each column represents a data point), target labels ${Y}\in\mathbb{R}^{1\times n}$, and weight vector $W\in\mathbb{R}^{d\times 1}$. Let $J(W)=\frac{1}{2}\|W^{T}X-{Y}\|^{2}$ be the Mean Squared Error (MSE) loss function, whose Hessian matrix with respect to $W$ is $H$. Suppose that $r$ principal components are used, $W_{r}$ is the matrix that consists of the selected eigenvectors of $X$, and let $Z=W_{r}^{T}X$. Then, the Hessian of $J(W_{r})$ with respect to $W_{r}$ has a condition number $\kappa(H_{r})$ that satisfies:
$\kappa(H_{r})\leq\kappa(H).$ (1)
Hence, in this case, PCsInit leads to a better-conditioned optimization problem, potentially resulting in more stable and faster convergence of gradient-based optimization algorithms compared to training with the original data. Now, recall that a function $f(x)$ is Lipschitz continuous if there exists a constant (Lipschitz constant) $L\geq 0$ such that for all $x,y$ in the domain of $f$: $||f(x)-f(y)||\leq L||x-y||$. $L$ represents the maximum rate at which the function can change. We have the following results regarding the Lipschitz constant for the first layer of PCsInit and PCsInit-Act:
Theorem 2.2
Let $x\in\mathbb{R}^{d}$ be the input vector, and let $W_{r}\in\mathbb{R}^{d\times r}$ be the weight matrix of the first layer, where $r\leq d$ is the number of principal components used. The columns of $W_{r}$ are the principal component vectors. Assuming that the bias term is zero, i.e., the output of the first layer is $\qquad h^{1}=W_{r}^{T}x$. Then, the Lipschitz constant for the first layer in PCsInit is:
$L_{1}=\sigma_{max}(W_{r})=||W_{r}||.$
Theorem 2.3
Consider the first layer of a neural network with the weight matrix $W_{r}\in\mathbb{R}^{d\times r}$ consists of the first $r$ principal components (PCs) of the input data, forming orthonormal rows, such that $W_{r}=Q$, where $Q\in\mathbb{R}^{r\times d}$ has orthonormal rows ($QQ^{T}=I_{r}$). Following this linear layer, an element-wise Lipschitz continuous activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ with Lipschitz constant $L_{\sigma}$ is applied. Then, the Lipschitz constant $L_{1}$ of the first layer operation $f_{1}(x)=\sigma(W_{r}^{T}x)$ with respect to the $L_{2}$ norm is $L_{1}=L_{\sigma}$.
Based on the Theorem 2.2, we see that the Lipschitz constant of the first layer in PCsInit is equal to the largest singular value of the weight matrix $W_{r}$, which is formed by the principal component vectors. This means that the maximum amount by which the output of the first layer can change for a unit change in the input is given by the largest singular value of $W_{r}$. Next, Theorem 2.3 shows that the Lipschitz constant of the first layer operation is simply the Lipschitz constant of the activation function used.
Now consider the Lipschitz constant of the first layer of a neural network based on the initialization method employed. PCsInit exhibits a fixed Lipschitz constant of $||W_{r}||$. Since $||W_{r}||$ is commonly orthonormal in implementation, this means the Lipschitz constant of PCsInit is a constant $||W_{r}||=1$, a characteristic it shares with orthogonal initialization mishkin2015all , which initializes weight matrices to be orthogonal and thus also possesses a constant Lipschitz constant. This inherent consistency can contribute to enhanced stability within the first layer. In contrast, He and Xavier initializations result in random Lipschitz constants that differ across initializations, introducing variability that can influence training dynamics.
2.2.2 Robustness to noisy data
One important point to bear in mind is that PCA decorrelates the input features. If the noise is independent across the original input features, then PCsInit also decorrelates the noise in the output of the first layer. Meanwhile, He and Xavier initialization can create correlations in the noise, as they involve random linear combinations of the input dimensions. In addition, we present the following statements regarding the robustness of PCsInit to noisy data, which illustrates that PCA projects the input noise onto the principal subspace. The variance of the noise in each principal component direction is scaled by the corresponding eigenvalue of $X^{T}X$.
Theorem 2.4
Assume that $\tilde{x}=x+\eta,$ where $\tilde{x}$ is the noisy input, $x$ is the clean input, and $\eta$ is the noise vector. In addition, assume also that $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, i.e., the noise follows a Gaussian distribution with zero mean and covariance matrix $\sigma^{2}I$. Here, $\sigma^{2}\in\mathbb{R}^{+}$ and $I$ is the identity matrix. Next, let the eigenvalues of $X^{T}X$ be $\lambda_{1},\lambda_{2},...,\lambda_{r}$. Then, for PCsInit, the noise propagated after the first layer is $W_{r}^{T}\eta$ follows Gaussian distribution with mean 0 and its covariance matrix is a diagonal matrix with entries $\sigma^{2}\lambda_{1},\sigma^{2}\lambda_{2},...,\sigma^{2}\lambda_{r}$.
Consider other initialization schemes for the first layer, if the input is noisy ($\tilde{x}=x+\eta$), the output of the first layer is $W_{1}\tilde{x}=W_{1}x+W_{1}\eta$. The noise component is $W_{1}\eta$. So, He and Xavier initializations transform noise by multiplying it with a randomly initialized matrix with a variance scaling. Meanwhile, orthogonal initialization rotates the noise vector without changing its norm. Next, we have the following theorem regarding the bound for the noise.
Theorem 2.5
Consider a neural network where the first layer’s weight matrix, $W_{r}\in\mathbb{R}^{d\times r}$, is initialized as in PCsInit, and therefore $W_{r}=Q$, where $Q\in\mathbb{R}^{d\times r}$ is orthonormal. Suppose that the input to the network is corrupted by additive white noise, i.e., $\tilde{x}=x+\eta$, where $x\in\mathbb{R}^{d}$ is the clean input signal and $\eta\sim\mathcal{N}(0,\sigma^{2}I)$ is the additive white noise. Then, the norm of noise component after the first layer is preserved, i.e., $||W_{r}^{T}\eta||=||\eta||.$
Also, let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$ be the neural network function, decomposed into layers $f=f_{L}\circ f_{L-1}\circ...\circ f_{1}$, where $f_{i}$ represents the $i$-th layer, $f(x)$ is the clean output, and $\tilde{f}(\tilde{x})$ is the noisy output. Further assume that each subsequent layer $f_{i}$ for $i=2,...,L$ is $L_{i}$-Lipschitz continuous. Then,
$||\tilde{f}(\tilde{x})-f(x)||\leq\left(\prod_{i=2}^{L}L_{i}\right)||\eta||.$
While the norm of the noise in the first layer is preserved for PCsInit, for other initialization methods (e.g., He, Xavier) where the first layer’s weight matrix $W^{\prime}$ does not have orthonormal rows, $||W^{\prime}\eta||$ is not guaranteed to be equal to $||\eta||$ and can be larger or smaller. In fact, we have $||W^{\prime}\eta||\leq||W^{\prime}||||\eta||,$ and it is possible that $||W^{\prime}||>1$. So, the noise can be amplified in the first layer. Next, we have the following result, which provide deeper details regarding the noise of layer 2 and subsequent layers:
Theorem 2.6
For any layer $\ell>1$, let $\rho^{\ell}$ is the activation function, $W^{\ell}$ is the weight matrix, and $b^{\ell}$ is the bias vector. Then, the output at layer $\ell$ is: $h^{\ell}=\rho^{\ell}(W^{\ell}h^{\ell-1}+b^{\ell}).$ Next, let $\eta^{\ell}$ represents the noise in the output of layer $\ell$ and assume also that $\rho^{\ell}$ is Lipschitz continuous with Lipschitz constant $L_{\ell}$, i.e. $||h^{\ell}-\tilde{h}^{\ell}||\leq L_{\ell}||W^{\ell}||\cdot||\eta^{\ell-1}||,$ where $\tilde{h}^{\ell}$ is the clean output and $h^{\ell}$ is the noisy output. Then, the bound for the noise propagated to the second layer specifically, which this isolates the influence of the first layer initialization of PCsInit compared to other techniques is
$\qquad||\eta^{2}||\leq L_{2}||W^{2}||\cdot||W^{1}||\cdot||\eta^{0}||.$ (2)
In addition, the general noise bound for any layer $\ell>1$ is:
$||\eta^{\ell}||\leq\left[\prod_{i=2}^{l}(L_{i}||W^{i}||)\right]\cdot||W^{1}||\cdot||\eta^{0}||.$ (3)
This theorem, which analyzes the propagation of noise through the first layer of PCsInit, suggests that PCsInit can handle noisy data in a way that allows for stable training. In addition, PCsInit, by capturing the essential structure of the data in the initialization, can act as a form of regularization. This regularization can prevent the network from learning spurious correlations caused by noise, improving generalization and robustness.
3 Experiments
3.1 Experiment Settings
To demonstrate the effectiveness of our proposed initialization strategies, we compare them with a standard Neural Network (NN) and PCA-NN on six big and small datasets whose details are found in Table 1. Among them, Parkinson and Micromass has number of features is high compared to the number of samples, MNIST deng2012mnist is image data, and HTAD garcia2021htad is a noisy sensor-collected dataset. For all datasets except MMIST, the experiments are repeated 10 times with 70%/30% train/test split ratio and the average is reported in the graphs. For the MNIST dataset, we simply used the already splitted train-test portions. Besides, Gaussian noise (mean 0, standard deviation 1) was added to each pixel feature after the initial standardization of the input data. More information can be found in the Appendix.
Table 1: Descriptions of datasets used in the experiments
@@ 3.1 Experiment Settings
Heart Dua:2019 $2$ $44$ $267$
Ionosphere Dua:2019 $2$ $32$ $351$
Parkinson Dua:2019 $2$ $754$ $756$
Micromass Dua:2019 $10$ $1087$ $360$
HTADgarcia2021htad $7$ $54$ $1386$
MNISTdeng2012mnist $10$ $784$ $60000$
For PCA, we choose the number of principal components so that a minimum of 95% of variance is retained, as commonly used in various research articles nguyen2022principle ; audigier2016multiple ; do2024blockwise ; nguyen2023principal . The width of the network used for each dataset has a width equal to the number of principal components retained. All models are trained with a total of 200 epochs using Cross Entropy Loss and Adam optimizer. The number of layers used for all datasets is 5. For PCsInit, the first layer weights were initialized using principal components derived from the training dataset, while the first layer of the standard NN was initialized using the specified initialization technique (He, Xavier, or Orthogonal initialization). Subsequent layers in both PCsInit and the standard NN, as well as all layers of the PCA-NN (which operated on pre-computed principal components), also employed one of three initialization techniques. For PCsInit-Sub, the initial PCA was performed on a randomly selected 20% subset of the training data. For PCsInit-Act, the ReLU activation function is used for the first layer. In the PCsInit family, the first layer is initialized with the principal components, and then it is frozen for the first 30 epochs. After that, we unfreeze this first layer and train the whole model for 170 more epochs. To facilitate a fair comparison, for all strategies, all layers except the first layers of the PCsInit family are initialized with the same weights. The first layer of PCsInit and its variants are initialized based on principal components. The experiments were run on a CPU of AMD Ryzen 3 3100 4-Core Processor, 16.0 GB RAM. The codes for the experiments is available at github.com/pthnhan/pcsinit.
3.2 Results and Analysis
@@
Figure 1: From left to right: result on the Heart dataset with He and Xavier initialization
@@
Figure 2: From left to right: results on the Ionosphere dataset with He and Xavier initialization
Note that in the captions of the resulting figures, when we say He initialization for example, we mean that all layers, except the first layers that initialized with a technique from the PCsInit family, are initialized with He initialization. The experiments results in figure 1-6 show the superior performance of PCsInit family compared to PCA-NN and NN in terms of esting accuracy and convergence speed, except for MNIST. Notably, the PCsInit-Act variation worked especially well when the He initialization was used for the layers after the first one. With this setup, it was one of the best methods for datasets with two classes (like Heart and Ionosphere). It also showed strong performance on larger, multi-class datasets such as MNIST and HTAD. The experimental results demonstrate that PCsInit variations and PCA-NN outperform the standard NN initialization across most datasets. For example, on the Heart and Ionosphere dataset (figures 1 and 2), the PCsInit family achieve clearly superior accuracy and convergence compared to PCA-NN and NN.
In addition, when dealing with noisy datasets, PCsInit family generally performed better than PCA-NN. For instance, on the HTAD dataset, a noisy data which is a common problem in sensor data, PCsInit and its variations were particularly effective (see figure 5). They were often the most accurate, learned quickly and steadily, and performed better than both PCA-NN and NN. This advantage was especially clear when He initialization is used for non-PCA layers in the PCsInit family. Similarly, on the MNIST dataset with added Gaussian noise(figure 6), the PCsInit-Act variation performed much like the other PCsInit variations. It started a bit slower but eventually achieved accuracy comparable to the normal NN. Meanwhile, the standard NN method started very well and learned quickly at first. In contrast, PCsInit and its variations began more slowly. However, they improved over time. By 200 training steps, their testing accuracy is similar to the standard NN. On the other hand, PCA-NN performed much worse on this MNIST data. This illustrate that the fine-tuning phase for the PCsInit family can provide significant room for improvement.
For datasets with large number of features compared to the sample size such as Micromass (figure 3) and Parkinson (figure 4), the PCsInit family gives pronounced improvement compare to NN and PCA-NN. On Micromass, while standard PCsInit still outperforms NN, its variations (PCsInit-Act and PCsInit-Sub) offer enhanced stability and competitive accuracy.
@@
Figure 3: From left to right: results on the Micromass dataset with He and Xavier initialization
@@
Figure 4: From left to right: results on the Parkinson dataset with He and Xavier initialization
Figure 5: From left to right: results on the HTAD dataset with He and Xavier initialization
Figure 6: From left to right: results on the MNIST dataset with He and Xavier initialization
4 Conclusion
In this work, we introduced PCsInit, a novel neural network initialization technique designed to embed Principal Component Analysis (PCA) within the network’s first layer. A key advantage of PCsInit, along with its variants PCsInit-Sub and PCsInit-Act, is the enhanced clarity and directness it brings to model explainability, avoiding the complexities associated with applying PCA as a preprocessing step. Across a variety of datasets, PCsInit demonstrates strong performance, often surpassing both PCA-NN and standard neural networks in accuracy and convergence speed. The theoretical properties of PCsInit, particularly its positive impact on Hessian conditioning, further solidify its value. Future research will delve into the broader applicability of PCsInit across different network architectures and data characteristics, including the exploration of Kernel PCA-based initialization.
@@ 4 Conclusion
In the future, we will examine the performance of the proposed approaches under various neural network architectures and various types of data (noisy, imbalanced). In addition, we will explore the potential usage of Kernel PCA instead of PCA for initializing neural networks, as well as analyzing deeper the choice of initialization techniques for the subsequent layers after the first layer for the PCsInit family.